There is an average of four accidents per year at a particular intersection.   What is the probability of more than one accident there next month? Hint: Use 1 month = 1/12 of a year to first get the number of accidents that are expected next month.       
Answer #1

Average rate = 4 per year

So, average rate = (4/12) per month = (1/3) per month.

This is Possion distribution because we have average value given.

Hence, probability of more than one accident in next month is 0.0446.

Answer is rounded to 4 decimal places.

Please comment if any doubt. Thank you.